### v2

# The Artificial Scientist: in-transit Machine Learning of Plasma Simulations

Abstract
Large-scale simulations or scientific experiments produce petabytes of data per run. This poses massive challenges for I/O and storage when scientific analysis workflows are run manually offline. Unsupervised deep learning-based techniques to extract patterns and non-linear relations from these large amounts of data provide a way to build scientific understanding from raw data, reducing the need for manual pre-selection of analysis steps, but require exascale compute and memory to process the full dataset available. In this paper, we demonstrate a heterogeneous streaming workflow in which plasma simulation data is streamed directly to a Machine Learning (ML) application training a model on the simulation data in-transit, completely circumventing the capacity-constrained filesystem bottleneck. This workflow employs openPMD to provide a high level interface to describe scientific data and also uses ADIOS2, to transfer volumes of data that exceed the capabilities of the filesystem. We employ experience replay to avoid catastrophic forgetting in learning from this non-steady state process in a continual manner and adapt it to improve model convergence while learning in-transit. As a proof-of-concept we approach the ill-posed inverse problem of predicting particle dynamics from radiation in a particle-in-cell (PIConGPU) simulation of the Kelvin-Helmholtz instability (KHI). We detail hardware-software co-design challenges as we scale PIConGPU to full Frontier, the Top-1 system as of June 2024 Top500 list.
Index Terms:
Exascale Computing, Computer architecture, High Performance Computing, Plasma simulations, In-memory computing, Data-driven modeling, Unsupervised learning, Dimensionality reductionI Introduction
HPC simulations capable of fully exploiting the compute power of supercomputers produce a wealth of high quality, complex data. Examples for such simulations include digital twins of the earth, whole device models of magnetic fusion reactors or simulations of advanced particle accelerators. The data rates and volumes of such simulations can be compared to instruments at large-scale research infrastructures such as tomography endstations at synchrotrons in photon science, high energy physics experiments, fusion research facilities or astro-particle observatories [1].
Exploiting such a huge volume of data optimally often focuses on fast data selection including data filtering, triggering, anomaly and outlier detection or intelligent data reduction. In-situ inference using pre-trained machine learning (ML) models has led to a broad inclusion of such techniques across scientific communities [2, 3, 4, 5].
Here we want to highlight an open challenge, namely the situation in which the stream of data produced by a simulation or any other source such as a scientific instrument remains at such a scale and quality that data reduction still demands in-transit analysis. In such cases, even if the volume of data is reduced to the minimum extent suitable for analysis, storage of such data for further offline analysis is impossible due to the physical constraints of the filesystem at hand.
The open challenge in our case is a particle-in-cell simulation of a Kelvin-Helmholtz instability (KHI) [8] as shown in Fig. 1 (more details in Section II). Instabilities are one of the most fundamental yet complex collective phenomena in plasma physics. Their growth rate, scale lengths and dynamics show a strong non-linear dependence on the plasma properties. Understanding and controlling instabilities is critical across all disciplines of plasma physics from magnetic and inertial fusion to space physics, plasma acceleration or astrophysics. Ultimately, reconstructing the complete phase space dynamics of a plasma from observations would provide us with the most detailed view of the growth and dynamics of instabilities and thus on the most central plasma properties at high temporal and spatial detail.
Reconstructing the complete phase space dynamics of a plasma requires processing of a wealth of information. One of the most challenging use cases in this class of problems is the in-transit training of data-driven complex machine learning models at extreme data rates, where training the model requires most of the data produced by the source. The model being trained itself is of such complexity that it can in principle incorporate the complexity of the total volume of data produced by the source, while at no time a total view of the data is available.
We note that a large number of advances in ML rely on iterative training [9], where the model repeatedly learns from the input data over multiple epochs, requiring permanent storage of the datasets. In the reference case studied in our work, this is not the case, as data is produced on demand and discarded after being used for training. We thus have chosen a continuous learning approach to address this challenge. Although methods for continuous learning exist [10], they have not had to consider the time constraints imposed by the speed of data generated from simulations [11]. Thus, adapting the continuous learning techniques to handle such constraints is a critical challenge in the field of ML for high rate streaming of scientific data.
The other challenge is to effectively navigate the memory hierarchy of heterogeneous computing systems, as it exposes enormous differences in throughput [12, 13]. While internal throughput in GPUs can easily reach the TB/s range on a single node, breaking down the throughput of massively parallel filesystems to the single node throughput reveals maximum performance of several tens of MB/s only. Each node often houses more than one GPU, hence available data storage and bandwidth demand data reduction, often done by discarding highly valuable data in practice.
We further note that the mismatch in throughput between memory hierarchies in modern HPC systems has been steadily growing, increasing the number of cases in which offline analysis is impossible and thus the urgency to find solutions for online analysis of data at high rates and volumes [12]. We argue that in case of such extreme rates online analysis of streaming data is the method of choice [14, 7], reducing the overall amount of data that needs to be stored while using computationally demanding algorithms to make optimum use of the data produced. However, this means that a complete view of all data produced is never available at any time of the analysis workflow and there is a need to preserve the extreme data throughput during the whole analytics pipeline.
In our work we showcase a real-world example and present a general solution that implements all relevant parts of online in-situ training of a large-scale model on the Frontier exascale supercomputer at ORNL [15]. We argue that the building blocks of our solution are general enough to be adapted to almost any use case of the type discussed above with only minor adjustments. We furthermore emphasize that besides presenting a full workflow, the additional unique value of this work is to highlight the multi-parameter optimization problem of fitting the available compute and data transfer resources to achieve the best result in the shortest amount of time. We remark that our technique is especially interesting in the case of irreversible data loss [16], and thus extends well beyond the case of large-scale simulations as high rate data sources presented here.
With a goal to automatically extract knowledge from large scale simulations, the paper makes the following contributions:
Computationally: We present a workflow that integrates a particle-in-cell simulation of the Kelvin-Helmholtz instability (KHI), supported by openPMD-streaming[14], with a PyTorch [17]-based ML application (which we refer to as MLapp from here on). We regard this orchestration as the Artificial Scientist. This orchestration, to the best of our knowledge, is the first of its kind.
Scientifically: We uncover correlations between emitted radiation and particle dynamics within the simulation in an unsupervised manner to create an inverse map from the observed radiation to the particle dynamics and to identify the aspects of the particle dynamics that are relevant to the remotely observable radiation signatures of KHI.
Using lessons learned from this orchestration and challenges faced, we provide input on what needs to be improved as we attempt to understand complex scientific data generated by instruments.
II The relativistic Kelvin-Helmholtz instability
The Kelvin-Helmholtz instability (KHI) is a well-known shear surface instability observed in plasmas [8]. It is driven by a self-amplifying cycle of small density or velocity fluctuations that lead to a growing magnetic field at the shear surface, which further amplifies the initial fluctuations as depicted in Fig. 1. This phenomenon occurs in astrophysical scenarios, including Saturn’s magnetopause [18], but also in fusion reactors and inertial confinement fusion (ICF) [19, 20]. While the theoretical understanding and simulation capabilities of KHI [21, 22] are well established, direct measurement of the predicted density changes remains a real challenge. Because of its well-understood dynamics, we chose KHI as a test case for our study to validate the results of our orchestrated workflow against existing knowledge.
A promising approach to the analysis of plasma dynamics is the study of the emitted radiation, which provides valuable information about the plasma properties and has been successfully applied in previous studies of the KHI [23]. However, this correlation process typically requires extensive post-processing and analysis, often spanning several years.
Here, we simulate the KHI using the plasma simulation code PIConGPU [24, 6] which utilizes the alpaka library [25] for performance portability and has been at the forefront of developing central technologies for GPU-accelerated plasma simulations. It was the first particle-in-cell (PIC) [26] code to run performant simulations on a variety of architectures in the world over many years [7].
III Reconstructing the local phase space dynamics of the relativistic KHI from observable radiation
In this section, we will showcase the challenges faced and solutions employed to solve an extreme-scale inverse problem via a complex HPC workflow coupling a GPU-accelerated plasma simulation (KHI/PIConGPU) to a large-scale in-situ neural network for in-situ learning. This workflow as a whole, is a novel approach, to the best of our knowledge.
Fig. 2 depicts the interrelations between observational data, such as radiation, and simulation state, such as particles, as both inputs and outputs to data-driven models. Our objective of approximating the inverse mapping from observed radiation to particle dynamics can abstractly be described by the information flow depicted in Fig. 2(a). For the model to remain tractable a reduced representation of the complex particle dynamics is required, which is learned by an auto-encoding model Fig. 2(b). To observe the extent to which this representation captures relevant features of the particle dynamics, it is worthwhile to train a surrogate model of the, computationally expensive, forward process of radiation emission (Fig. 2(c)). These three tasks integrate into our ML model, described later, also in Fig. 7.
It was shown in 2013 [6] that the computation of high quality, high resolution observational data such as radiation spectra can only be performed in-situ, as storing the ground truth data, the phase space trajectory data of each particle in the simulation at each time step, is impossible due to the aforementioned throughput hierarchy. As an example, scaling to a moderate 25 % of the Frontier system, one would be faced with 1 PB of data for every time step. With total time steps per simulation of the order of a few thousand time steps and time step durations of the order of 0.1 to 1 s, we frequently observe data rates of 1 to 10 PB/s for particle trajectory data and theoretically require total volumes on disk of the order of 10 EB. Depleting storage in as little as . We thus end up with the conundrum to employ data-driven models for solving a large-scale ill-posed inverse problem.
III-A A complex in-memory workflow to reconstruct local phase space dynamics from radiation
A particular challenge for the reconstruction of local phase space dynamics is that their radiation signatures are highly convoluted, composite quantities that combine multiple facets of the ground truth which to restore. This observation guides our approach at learning from a live PIC simulation, as the posed problem is now diagnosed as ill-posed. In detail, the observed radiation spectrum is both spectrally and angularly resolved and the observed intensity at each wavelength and direction varies between time steps according to the changes of the phase space dynamics. The observed intensity in each direction and at each wavelength is a weighted sum of all contributions from all particles.
Hence here are the steps to be taken:
Simulate a sufficiently large domain at high spatial and temporal resolution to capture both the local plasma dynamics as well as the collective contributions of different plasma regions to the total radiation spectrum.
At each PIConGPU time step:
Calculate the radiation spectrum directly in-situ as the sum of the contributions from all radiating particles in the simulation at each time step, with high angular and spectral resolution and bandwidth.
Collect the particle phase space data, positions and momenta, and the complex amplitude and phase for each wavelength and direction of the radiation.
Prepare the collected data for an ML model by finding a suitable encodings for spectral and phase space data.
Stream that data to an ML model and asynchronously train the model.
Stop the simulation after observing the complete physical process under investigation.
III-B The technical challenge: Exascale data streaming and continual learning
An important design goal for the implementation of the above data processing scheme is to write no simulation data to the filesystem during the entire workflow. File I/O can certainly be initiated when desired, but is independent of and not necessary for the workflow.
While in-situ approaches have been previously used to compute scientifically relevant analyses through plugins within the same application context [6], this is no longer possible for the envisioned ML analysis, as the vastly incompatible software stacks for the C++ simulation code PIConGPU and the PyTorch-based MLapp as well as their different scaling properties inhibit such an integration into a strongly-coupled monolithic application framework. Hence, simulation and the MLapp are modularized into two separate applications that communicate through a standardized data interface, a pattern that we refer to as loose coupling, since the standardized data interface allows to plug together a data processing pipeline more flexibly, i.e. loosely than before. When the implementation for such workflows avoids writing intermediate data to disk, we refer to them as in-memory. In-memory workflows include such setups where data is not moved, staying within one application’s memory (in-situ), and alternatively setups that move data from one into another application’s memory (in-transit) (Fig. 3(a)). The latter becomes necessary when transitioning from file-based workflows to in-memory workflows without sacrificing flexibility in the data exchange pattern, as well as in response to codes that scale differently.
The whole simulation, all data produced by the simulation and ingested by the model, and the whole model all reside in-memory and data is distributed and optionally reduced in-transit (Fig. 3(b)).
The physical workflow outlined in III-A can be directly translated into a technical setup that is the main subject of this work:
Start a scalable, GPU-accelerated PIConGPU simulation of sufficient size and resolution, utilizing a large subset of nodes on Frontier.
Schedule PyTorch and N/RCCL alongside the PIConGPU simulation. (N/RCCL - NVIDIA and ROCm collective communication library)
Optimize scheduling for data transport and locality, physical system size and resolution, simulation and training time, total time of solution and use of resources (options in Fig. 3(c)).
At each simulation time step
stream particle momenta and positions and spectral data to a large-scale ML model,
transform the phase space and spectral data from its simulation format to an optimum format for the ML mode and
train the model concurrently.
Repeat the workflow for a sufficient number of time steps to cover all relevant stages of the plasma instability simulation with respect to the dynamical evolution and amount of training data needed.
For supporting an online workflow which stores no intermediate data such as phase space trajectory and radiation data to disk, the model must be continuously trained online from the series of time step data. This requires the model to be trained by subsequent snapshots of the evolution of plasma and radiation data. For this, continual learning is the methodology of choice as outlined in Section IV-C.
We strongly believe that a large class of inversion problems for complex, large-scale systems that produce data streams face similar challenges and that our solution presented here for the first time introduces all relevant building blocks and technologies to address these challenges.
IV Methodology
IV-A Particle-in-cell simulation and its setup
To simulate KHI, we employ the fully kinetic 3D3V relativistic particle-in-cell (PIC) method [26] using the open-source code PIConGPU [6, 24]. The PIC method iteratively models plasma dynamics by advancing particle positions and velocities in response to electromagnetic forces, while the electromagnetic fields are updated based on the electric current caused by the particle motion.
For modeling the KHI with PIConGPU, the smallest volume we simulate is cells on 16 AMD MI250X GPUs using cubic cells of and a time step duration of at a particle density of . We initialize two counterpropagating plasma streams (as shown in Fig. 1) with a normalized velocity of and initialize 9 particles per cell.
To prove that PIConGPU can scale on a large system for general plasma physics cases, we use a more challenging test case [27] than the KHI as a scaling benchmark, with a higher particle-per-cell ratio. Fig. 4 shows scaling runs on ORNL’s Summit and Frontier.
On Frontier with a total of AMD MI250X GPUs across nodes [28], a scaling of PIConGPU’s Figure of Merit (FOM), i. e. the weighted sum of the total number of particle updates per second (90 %) and the number of cell updates per second (10 %), is demonstrated as shown in the weak scaling results. In this particular configuration, the simulation included macroparticles in grid cells, compared to our base FOM run on Summit in 2019, where only macroparticles in cells could be used.
On Frontier, one thousand time-steps completed in a mere 6.5 minutes. This is due to several factors. Spatial domain decomposition distributes computational domains across GPUs and computers, exploiting data locality and enabling parallelization. Efficient data structures, such as supercells in PIConGPU [29], optimize data access patterns within these domains, while asynchronous communication strategies between compute nodes minimize communication overhead when exchanging information between different domains. In addition, usage of the libraries PMacc [30] and alpaka [25] abstract data management and hardware-specific optimizations, enabling portability across architectures.
PIConGPU uses openPMD [31, 32] to output data. By using openPMD, we stream particle data in-transit via ADIOS2 [33] (see Section IV-B) directly to the ML framework, thus avoiding filesystem limitations [12]. In addition, PIConGPU supports strongly-coupled in-situ processing and analysis methods, among which we use the far-field radiation plugin for our study. In contrast to the intrinsic field solver, the far-field radiation plugin [34] calculates radiation emissions using the Liénard-Wiechert potential approach [35], thus overcoming common limitation of the PIC algorithm, as it resolves much higher frequencies, directly predict observable spectra, and correctly quantify coherent and incoherent radiation [36]. However, it comes at a high computational cost, scaling with the number of particles, frequencies, and directions, easily exceeding the computational limits of the underlying particle-in-cell simulation [37].
Since the computational cost of the far-field radiation plugin can be significant, forward calculations can become prohibitively expensive for a wide range of simulation parameters. Furthermore, inversion from observed radiation back to particle dynamics remains a challenging task, requiring ML-based innovative approaches. Thus, we implemented a novel ML approach (see Section IV-C) that was trained on both particle distributions and their associated radiation using streamed data from PIConGPU. After training, the ML model should generate particle distributions based on the input spectra, providing a potential solution to the inversion problem and improving our ability to interpret plasma dynamics based on observable radiation.
IV-B Data streaming via openPMD
Considering that PIConGPU as well as the MLapp both rely significantly on very different software stacks, launching both parts of the setup in the same application context is not feasible. As detailed previously in III-B, we follow a loose-coupling rather than a tight-coupling approach [14], launching both applications separately. Output created by PIConGPU must be moved across the application boundary to the MLapp, without impacting the scalability of the setup through use of limited resources such as the parallel filesystem. The solution pursued in this paper is an in-transit workflow that keeps data in memory at all times and writes no intermediate data to disk.
The in-transit I/O solution is put in place by openPMD[31] implemented within PIConGPU and the MLapp. openPMD [32] is a data standard for particle-mesh data built for F.A.I.R. scientific I/O in HPC software that needs to scale, and it is employed in a diverse ecosystem of simulation, analysis and visualization codes[38]. In turn, the I/O building block for the orchestration of the Artificial Scientistis modular and generically reusable for other setups of this kind.
As an open data format, openPMD can be implemented in various file formats such as JSON, TOML, HDF5 and ADIOS2, all supported by the reference implementation openPMD-api, of which we further consider ADIOS2 in this study due to its support for in-transit I/O in HPC settings, as shown in Fig. 5, This flexible choice of backends for scalable and quickly adaptable I/O solutions gains outstanding relevance on exascale like Frontier where even modern solutions for classical I/O via parallel filesystems, e. g. the provisioned Orion filesystem, cannot be expected to scale up to full system size as a base for a loosely coupled data analysis pipeline.
The in-transit data strategy is built upon the Sustainable Staging Transport (SST) [39, 40] data engine of ADIOS2 that connects one parallel data producer code to an arbitrary number of parallel data consumer codes, hereby opening connections only between processes that actually share data. For adapting to different system architectures, the SST engine implements different network transport technologies (data planes), including TCP (non-scalable fallback), libfabric, ucx and the MPI_Open_port() API of MPI. For this benchmark, we consider the low-level libfabric data plane on top of the system’s CXI provider [41] which allows, yet also requires more manual finetuning, as well the MPI data plane which leverages the MPI implementation’s performance tuning.
The use of HPC interconnects such as the Slingshot network on Frontier or Infiniband on e. g. Summit is an important cornerstone in the general applicability of this I/O building block, since this I/O approach thus retains the flexibility of setups based on the parallel filesystem. In particular, there is no lock-in to a specific pattern for data exchange, compared to approaches based on e. g. local SSDs or shared memory. While local data exchange within a node remains preferred for scalability, this I/O approach naturally extends towards patterns such as staging within a neighborhood of nodes (for scheduling reasons or for implicit load balancing via streaming) or a fan-in pattern (for data reduction purposes), both of which are potential directions to pursue.
Before applying this I/O solution to the MLapp, we demonstrate its scalability to the full system using a synthetic benchmark that runs PIConGPU’s KHI scenario full-scale and streams its particle data into a synthetic no-op consumer that performs no computation beside measuring the performance of this I/O operation and only discards received data. Employing the no-op consumer gives us a testbed for full-system scaling runs of a particle data stream fed by PIConGPU, helping us identify and eliminate scaling issues before applying the full PIConGPU+MLapp pipeline. In that context, the major challenge is to scale to a large number of parallel instances, while the amount of data streamed per instance comes as a secondary scaling target.
The PIConGPU KHI simulation is run from half (4096 compute nodes) to full-scale (9126 compute nodes) of Frontier, producing of particle data per compute node and time step. Fig. 3(c) (bottom left) shows the processes running within each compute node and the data flow between them. At each scaling run (weak scaling), five time steps are sent from PIConGPU to the no-op consumer, which then measures the time needed for loading the data. The perceived parallel throughput is computed based on this measured time and the global data size, i. e. the previously mentioned per node multiplied with the number of nodes. While including communication overhead, this perceived parallel throughput has been shown in [40] to be a close approximation for the real throughput in this kind of setup.
The benchmarks in Fig. 6 demonstrate the feasibility of this I/O strategy for full-scale workflows. Depicted are boxplots of all single measurements. Most notably, we observe a maximum perceived parallel throughput of 20 - 30 TB/s which compares outstandingly against the 10 TB/s bandwidth of the parallel Orion filesystem, a scaling limit which we circumvent and exceed by not using the filesystem for intermediate data. Furthermore, these results can also compare well against the 35 TB/s aggregate write bandwidth of the SSDs installed locally in the compute nodes [42].
For the libfabric-based benchmarks, we initially used an implementation that enqueued all read operations to the network at once and waited for replies (labeled in Fig. 6). While these results signify the best achieved per-node throughput of 3.5 - 4.7 GB/s at 4096 compute nodes (a total of 14.1 - 18.8 TB/s), it turned out that this strategy did not scale to the full system, so we employed an alternative that enqueued operations in batches of 10 operations. While this turned out to scale to the full system, it came at a notable cost in performance, visible in a per-node throughput of 1.9 - 2.6 GB/s at 9126 compute nodes (a total of 16.5 - 23.0 TB/s). Conversely, the MPI data plane yields a per-node throughput from 2.6 - 3.7 GB/s at 4096 compute nodes (a total of 10.5 - 14.9 TB/s) to 2.4 - 3.3 GB/s at 9126 nodes (a total of 21.4 - 29.5 TB/s), the best achieved system-wide parallel throughput from this experiment.
We conclude that the MPI data plane brings default good performance, while the libfabric data plane’s lower-level control can bring performance improvements but at the cost of necessary fine-tuning beyond this study’s scope. Since only a single instance of the reader code was running per node, its throughput ( - , across all cases) compares against the max possible throughput of a single HPE Slingshot NIC at . This implies that further speedup can be achieved by parallelizing the reader in the actual PIConGPU+MLapp pipeline. The benchmarks suggest that the MPI data plane be used, which can be explained by leveraging system-specific performance tuning in the MPI implementation. All regular measurements range between - .
IV-C Machine learning
The task of in-transit ML is to extract patterns from the live observations and thereby learn correlations between variables and reduced representations. Here, we focus on predicting particle dynamics from radiation spectra, which constitutes an ill-posed problem in that any given radiation spectrum can potentially be produced by various types of dynamics. When learning in-transit from a simulation, this problem is reduced from an intractable general case to the question of: which type of local particle dynamics under the given physical setup, i. e. the relativistic KHI, gave rise to an observed radiation spectrum observed in one direction, which, however, remains an ill-posed problem. Many ML techniques, especially feed-forward networks, tend to predict the mean of possible inversions, even if this itself is not a valid solution [43]. One class of techniques specifically designed to learn by sampling from conditional multi-modal distributions, such as those that occur as solutions to our ill-posed problem, are flow-based models [44, 45] and invertible neural networks (INNs) [46]. These learn an invertible transformation, or change-of-variables, between a chosen posterior distribution, often a normal distribution , and a prior described by the training data, in our case the distribution of observed particle dynamics . One specific difference between them is, that INNs also learn the deterministic forward function defining the inverse problem, in our case the mapping between samples of particle dynamics and radiation spectrum . Put differently, they also learn to predict the condition under which a given sample of the prior may occur.
Particle configurations in phase space of the plasma dynamics simulated here can be very complex, such that a small sample of particles cannot adequately represent the dynamics. For INNs and flow-based models this poses a challenge, because the change-of-variables approach [47, 48] dictates, that the information volume, i. e. the number of activations at the input and output of each coupling block, must be constant throughout the network, which results in a very large network when working with a prior distribution of dimensionality on the order of . In this setting, the number of network parameters can be kept in check by using affine flows [49, 44] with designed convolutional sub-networks which should also be invariant to transposition of particles in the input vector to be efficiently trainable.
When generating point cloud (PC), flow-based models are typically tasked with sampling them point-by-point, i.e. the model learns a probability density that describes the particle density in phase space. The main drawback here is, that the model has to learn the many-modal distributions which can occur in the KHI phase space on a single-particle basis and can not longer generate multiple particle distribution for a single condition, i.e. not distinctly produce all PCs solving an ill-posed problem. Another issue is, that, unless we use an encoder (conditioning network) producing very small latent vectors of around six dimensions for the radiation, the dimension of the condition is much larger than the dimension of the prior while the network has to be sampled times to retrieve a useful particle distribution.
To overcome the above inefficiencies, we opted to view the particle configuration as a point cloud (PC) and reduce the dimensionality of the prior by tasking an outer network to learn latent representations of the phase space, which are also invariant with respect to transposition of particles. This outer network can in principle be an autoencoder (AE) whose latent representation is passed to the INN as the prior . In practice, the INN will not generate exact latent vectors on its backward pass, but variations, which will cause the AE’s decoder to fail at reconstruction. Therefore, we employ a variational auto encoder (VAE) instead, which, albeit harder to train, intrinsically trains its decoder to be robust against the variations produced by the INN.
As depicted in Fig. 7, there are three main blocks of the network architecture:
The encoder (light green) follows an architecture proposed in [50], which by itself is a simplified version of PointNet [51], using convolutions acting on the coordinates of particles separately to gradually (channels: ) extract 608 features per particle, followed by a max-pooling to obtain a transposition-invariant feature set, which is then processed by two multi-layer perceptrons (MLP), each with hidden features, to obtain predictions for the mean and standard deviation to sample the 544-dimensional latent vector .
The decoder (cyan) uses a single fully-connected layer (, reshaped to to transform the latent before using 3D deconvolutions (channels: , kernelsize: , stride ) to upsample to particles. While generating a larger PC that matches the input size of particles would be beneficial to obtain a smoother loss surface, training a network that generates only particles as output is also feasible and will reduce the VAE network size. In the present case, using particles as an input during training, the network can learn to generate important particles to represent the target PC. Randomly sampling fewer particles from the input would not yield a sufficient representation.
The INN for the inverse mapping (violet) is built from four Glow coupling blocks [52] using MLPs with hidden layers as subnets. The hidden layer sizes were chosen to form a bottleneck and reduce to a computationally efficient power of two.
The loss function to be minimized during training of this architecture consists of five terms, two connected to the VAE and three for the INN, detailed below:
As the reconstruction loss of the VAE, we use Chamfer’s distance (CD) [53], which is the most commonly used differentiable function to compare PCs in literature, because it is efficient to compute.
Regularization of the VAE’s latent space is controlled by Kullback-Leibler divergence (KL) [54].
The resulting spectrum of the INNs forward function is compared to the ground truth using mean-squared error
The posterior normal distribution, produced by the INN’s forward pass, and the prior distribution, produced by the INN’s backward pass, are both evaluated using maximum mean discrepancy (MMD) using an inverse multi-quadratic kernel, as proposed in [46].
The total loss corresponds to the weighted sum
The numerical values for the coefficients were empirically tuned to optimize performance of the AE reconstruction and the INN forward and backward predictions.
While training on the continuous data stream of non-steady configurations, we employ experience replay (EP) [55] to avoid catastrophic forgetting of earlier simulation time steps while training on later ones. To integrate with the streaming setup, we implemented this approach to continual learning as a separate component, called training buffer, between the streaming receiver and the training loop. This component holds two buffers: a now-buffer, which holds the latest samples received. Each time new samples are received, they are prepended to the now-buffer and any element which would then exceed the buffer size is removed and added to the EP buffer. The EP buffer holds a maximum of elements. When new samples are to be added to the EP buffer when it is full, a randomly chosen item is removed. For a single iteration of the training loop, a batch of data is requested from the training buffer. This batch constitutes the training batch. It is generated by taking random samples form the now-buffer and random samples from the EP buffer, yielding a training batch size of . In order to increase training success, we perform iterations of the training loop per single time step from the data stream. For each iteration, a new training batch is generated from the now- and EP buffer.
Separating the EP schedule from the training loop via our training buffer allows us to control how many batches we iterate per sample time-step produced, as long as we have some leeway to stall the running simulation if need be. This is crucial to allow the optimizer some amount of exploration, which can only happen sequentially and hence a smaller number of training iterations cannot be compensated by the large batch sizes of data-parallel training.
For all training runs, we use the Adam optimizer [56] with , , and an weight decay strength of . For data parallel training we employ PyTorch’s Distributed Data Parallel (DDP) module with its N/RCCL backend. Learning rates are scaled following a square-root rule [57]. In order to accelerate learning with large batch sizes, we train layers belonging to the VAE at a learning rate higher by a factor than that of the INN.
Technical Challenges: The main drawback of CD, used as the VAE’s reconstruction loss, is its lacking sensitivity to point density. It also fails to distinguish some point configurations [58]. Both of these problems are overcome by the earth mover’s distance (EMD) [59] loss, but at much higher computational cost111We observed about a increase in batch times when using EMD [60] vs. a simple implementation of CD.. Still, a rather compute and memory efficient differentiable implementation of EMD is available in the geomloss python package [60], which, instead of relying entirely on PyTorch, internally uses the KeOps library [61, 62] for performance. Unfortunately, the KeOps library is implemented directly in CUDA, and thus cannot be used on AMD GPUs in Frontier. Hence, we were barred from employing EMD loss in our runs on Frontier. Perhaps the community needs a HIP version of the KeOps library.
IV-D Setup on ORNL Frontier system
While distributing nodes exclusively to either PIConGPU or to the MLapp, inter-node setup in Fig. 3(c), is easier to achieve in Slurm (and this strategy would be supported by our flexible I/O solution), we decide to use the intra-node setup in Fig. 3(c) and share each compute node between PIConGPU and the MLapp. This has the advantage that data exchange mostly needs not leave the node, but requires a more specific resource allocation: Each single compute node no longer runs homogeneous parallel sub-jobs, but has to locally assign resources to heterogeneous applications. For our setup, this local assignment gives 4 GCDs to PIConGPU and 4 GCDs to MLapp. In openPMD and ADIOS2, each reader application decides on its own which remote datasets to load. Here, the distribution of loaded data regions (i. e. local blocks within the global dataset) is configured such that data is shared within node boundaries.
Due to the plugin-based structure of PIConGPU, both the particle and radiation input required by MLapp are provided by distinct output plugins of PIConGPU, meaning that two parallel data streams are opened between PIConGPU and the MLapp. Each plugin prepares the current time step’s data and, upon having all data ready, ADIOS2 will gather the metadata, including remote read addresses, to parallel rank 0 and provide the time step to the reader. The reader performs remote reads to its liking and closes the time step, indicating to the writer that the data can now be dropped.
A unique and important requirement of the presented workflow is the ability to not only run but also scale on a large number of Frontier’s AMD MI250X GPUs. Our machine learning model is small enough to fit on a single GCD; therefore, parallel training of this model is done using data parallelism where copies of the model are distributed across GCDs with each copy of the model receiving different chunks of data to train on. Once each model computes its gradients, all the instances of the model must do a collective all-reduce communication to average the gradients. We expect the scaling of the workflow to critically depend on the optimization of this all-to-all communication in PyTorch DDP, since communication within the PIConGPU simulation is only between next neighbors. We experienced a hard limit to scaling in our current setup: The all-to-all communication between PyTorch DDP ranks using the N/RCCL backend hits system limitations on the possible number of open sockets beyond 100 nodes. Potential options to circumvent this limit include using a libfabric backend for N/RCCL or using PyTorch DDP’s MPI backend. While the N/RCCL backend was able to support PyTorch up to the scale used in this study, we can already conclude that further extensive evaluation of the communication backends within PyTorch will become mandatory to provide a stable base for our near future scaling studies on Frontier. This area of research appears to be still under development, to the best of our knowledge [63].
V Results
V-A Weak-Scaling
When training any kind of statistical model, in our case a deep neural network, one has to ensure that for each characteristic of the domain one hopes to learn, there are sufficient examples available in the dataset. When the dataset is provided in-transit, redundancy must still be built into the setup, either by using a large simulation volume or running multiple independent simulations concurrently, to allow the network to learn repeated patterns of the simulations and thereby distinguish relevant domain features, such as physical relationships, from noise. For this reason, in this work, we focus on a weak-scaling study of the proposed pipeline, as we anticipate this to be the most likely use-case. To evaluate parallel scaling as well as learning scaling, we performed simulations with in-transit training from 8 up to 96 nodes. The lower limit of 8 nodes is set by physical size constraints of the simulation. Fig. 8 shows the weak scaling behavior of the pipeline on Frontier. We define efficiency as the ratio of runtime at a specific size over runtime of the smallest size. Before averaging the times per training batch, we removed outliers more than four standard-deviations from the mean, because on Frontier we observed single batches taking more than the mean time.
While ideal scaling should correspond to 100 % efficiency across all sizes, communication overheads reduce efficiency in practice. Here, the efficiency of in-transit training reaches around 35 % at 96 nodes. We attribute the low parallel efficiency to two main reasons: One of them is the inevitable all-to-all communication between PyTorch ranks that takes place to average the gradients during each backward pass and obtain a synchronized result on all ranks, which accounts for a deficit of , which is still a respectable result achieved by overlapping computations and communication in PyTorch DDP. The other is the computation of the two MMD loss terms, which amount to matrix dot products with data distributed across all ranks. Due to torch not offering a distributed primitive and the matrices involved being relatively small, even at the largest achieved scale, we opted for a naive implementation which replicates work between nodes. The global communication required here222torch.distributed.all_gather_into_tensor breaks the torch computational graph, i.e. synchronizes graph execution with host code at the invocation site. To our knowledge there is no high-level API available to stitch these operations into the graph to allow them to be executed asynchronously to both the host and each other, i.e no enqueueing of host functions or events.
V-A1 Scaling Model Training
The performed scaling runs, training with a batch size of per GCD, correspond to total batch sizes of 256 to 3072 on 32 to 384 GCDs, respectively.
In our scaling runs, we explored training from scratch at large scale, sampling up to batches per sample consumed from the stream and found learning success up to about . We settled on a base-learning rate of , scaled according to the square-root scaling rule [57] depending on the actual number of GCDs. We observe a training speed up with larger batch sizes. That is, the increased learning rate overcompensates the training slow down due to larger batch sizes. Furthermore, it became evident that the three losses connected to the INN converged best at lower learning rates, while the VAE only found better minima at the highest learning rate, showing that separate learning rates and of the encoder/decoder block and the inversion block, respectively, need to be applied at large scales.
Within the scope of this study, we have to conclude, that for in-transit training at very large batch sizes, hyper parameter studies do not transfer from small-scale experiments and thus have to be performed at scale. To enable in-transit learning at the scales we tested or beyond, comprehensive studies of the relations between the block learning rates and , batch sizes, and maybe even loss weights in Eq. (1) have to be performed. Exploring this parameter space for the large scale training requires the full study to be systematically performed at these scales (up to 384 GCDs), either by employing a streamed simulation or the same training data stored on disk. We argue, that the latter is challenging at scale due to the lack of disk capacity and bandwidth.
V-B Quantifying the predictive capabilities of the machine learning method
To evaluate the performance of the trained model, we invert the radiation spectra back to the original momentum distribution, focusing on the momentum component in the following discussion (see Fig. 9). As this inversion is an unsolved problem in physics, we do not expect good agreement, but would be satisfied with partial reconstruction of the distribution.
Given spectra from the bulk of the plasma, the model reconstructs the momentum distribution quite well (blue and red Fig. 9(c)). The mean of the predicted momentum agrees with the distribution from PIConGPU (Fig. 9(b)). We also observe that the network learned a fundamental aspect of special relativity: the Doppler shift, to distinguish between plasma streams approaching and receding from the detector, since it reproduces the correct cutoff frequency when predicting spectra from particle data (Fig. 9(a) dotted lines).
The momentum prediction from the KHI shear surface was expected to be more difficult since particles from both directions cross in the shear surface, resulting in two separate and distinct particle populations. The ML reconstruction consistently predicts these two distinct populations. This is consistent with the interpretation, that the INN has learned unsupervised that the vortex region undergoes more acceleration and thus emits more intense radiation than the bulk of the plasma. However, the momentum prediction is less accurate: the mean of (one of) these populations often does not match the PIConGPU data (Fig. 9(b vs. c) green). We suspect that this is partly due to our choice of region boundaries on the shear surface, which results in a much lower sampling rate of one flow direction compared to the other, leading to a lower accuracy in predicting the mean of the lower peak of the momentum distribution. It is also possible, that this problem arises from the VAE’s decoder having difficulties reconstructing discontinuous distributions. Nevertheless, the successful prediction of shear regions indicates that the encoder extracts crucial features from the input data to recognize whether a region contains a vortex and encodes them in latent space.
Despite suboptimal reconstruction of vortex momenta, the ML prediction still allows to clearly identify KHI regions. Previous work on PC generation [64] observed that the latent space of an AE may contain more information than the decoder can reconstruct with satisfactory fidelity and that more sophisticated generative models, such as GANs or flows, can learn to produce more detailed reconstructions when conditioned on the same latent space.
While the prediction accuracy for the inversion from radiation to momentum is still limited, our workflow shows promising potential for automated learning of physical relationships from in-transit data of large-scale simulations and the identification of regions of instability is already a major step forward. This identification by the network may be obvious to a domain expert, but the Artificial Scientist learned the distinction unsupervised. The work has demonstrated, as a proof of principle, that it is possible to learn correlations from a physics simulation on-the-fly. This is a promising first result towards large-scale in-transit learning for a non-steady state processes.
Future work will explore more sophisticated network components such as (a) Deeper INNs with more powerful subnet-architecutres [44], (b) a wealth of more sophisticated PC decoder networks to generate higher-fidelity depictions of predicted particle distributions [64], and (c) more complex encoders architectures [51], ideally bringing contrastive learning approaches [65] to point clouds to learn better latent representations.
VI Conclusion
By training a machine learning model on-the-fly on a PIConGPU simulation of the Kelvin-Helmholtz instability, where data is streamed via openPMD with ADIOS2, we demonstrate a realization of the Artificial Scientist solving an ill-posed inverse problem from plasma physics. This allows us to circumvent the lack of adequate disk capacity and bandwidth, often a bottleneck in large-scale systems, by enabling data to reside in-memory and be distributed via network between the nodes. Scientifically, this work also demonstrates as a proof of principle, that it is possible to learn correlations from a physics simulation on-the-fly. On full Frontier, we demonstrate streaming at scale using openPMD as well as performance studies at scale for general plasma physics.
Acknowledgments
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
The research was supported by Laserlab-Europe EU-H2020 GA no. 871124.
The support work from the ADIOS team was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program, under the “RAPIDS Institute”.

### v3

@@
 We employ experience replay to avoid catastrophic forgetting in learning from this non-steady state process in a continual manner and adapt it to improve model convergence while learning in-transit.
-As a proof-of-concept we approach the ill-posed inverse problem of predicting particle dynamics from radiation in a particle-in-cell (PIConGPU) simulation of the Kelvin-Helmholtz instability (KHI).
+As a proof-of-concept, we approach the ill-posed inverse problem of predicting particle dynamics from radiation in a particle-in-cell (PIConGPU) simulation of the Kelvin-Helmholtz instability (KHI).
 We detail hardware-software co-design challenges as we scale PIConGPU to full Frontier, the Top-1 system as of June 2024 Top500 list.
@@
 In-situ inference using pre-trained machine learning (ML) models has led to a broad inclusion of such techniques across scientific communities [2, 3, 4, 5].
-Here we want to highlight an open challenge, namely the situation in which the stream of data produced by a simulation or any other source such as a scientific instrument remains at such a scale and quality that data reduction still demands in-transit analysis.
-In such cases, even if the volume of data is reduced to the minimum extent suitable for analysis, storage of such data for further offline analysis is impossible due to the physical constraints of the filesystem at hand.
-The open challenge in our case is a particle-in-cell simulation of a Kelvin-Helmholtz instability (KHI) [8] as shown in Fig.
+Here we want to highlight an open challenge, namely the situation where a stream of data produced by a source – like a large-scale simulation or a high-resolution and high-repetition-rate experiment – is of such a scale and quality that, even after simple data reduction, storage for further offline analysis is impossible due to the physical bandwidth or capacity constraints of the filesystem at hand.
+In such cases data reduction already demands complex in-transit analysis.
+We illustrate this challenge with a particle-in-cell simulation of a Kelvin-Helmholtz instability (KHI) [8] as shown in Fig.
 1 (more details in Section II).
-Instabilities are one of the most fundamental yet complex collective phenomena in plasma physics.
-Their growth rate, scale lengths and dynamics show a strong non-linear dependence on the plasma properties.
-Understanding and controlling instabilities is critical across all disciplines of plasma physics from magnetic and inertial fusion to space physics, plasma acceleration or astrophysics.
-Ultimately, reconstructing the complete phase space dynamics of a plasma from observations would provide us with the most detailed view of the growth and dynamics of instabilities and thus on the most central plasma properties at high temporal and spatial detail.
+Instabilities are one of the most fundamental, yet complex collective phenomena in plasma physics.
+Understanding and controlling them is critical across all disciplines of plasma physics from magnetic and inertial fusion to space physics, plasma acceleration or astrophysics.
+Instability growth rate, length scales and dynamics show a strong non-linear dependence on the plasma properties.
+Reconstructing the complete phase space dynamics of a plasma from observations would provide the most detailed view of the growth and dynamics of instabilities, and thus of the most central plasma properties at high temporal and spatial detail.
 Reconstructing the complete phase space dynamics of a plasma requires processing of a wealth of information.
-One of the most challenging use cases in this class of problems is the in-transit training of data-driven complex machine learning models at extreme data rates, where training the model requires most of the data produced by the source.
-The model being trained itself is of such complexity that it can in principle incorporate the complexity of the total volume of data produced by the source, while at no time a total view of the data is available.
+One of the most challenging use cases in this class of problems is the in-transit training of data-driven complex ML models at extreme data rates, where training the model requires most of the data produced by the source.
+The model being trained itself must be of sufficient complexity that it can, in principle, incorporate the complexity of the total volume of data produced by the source, while at no time a total view of the data is available.
 We note that a large number of advances in ML rely on iterative training [9], where the model repeatedly learns from the input data over multiple epochs, requiring permanent storage of the datasets.
 In the reference case studied in our work, this is not the case, as data is produced on demand and discarded after being used for training.
+Due to the size of data used for training and memory constraints, no total view of the training data is available at any given time.
 We thus have chosen a continuous learning approach to address this challenge.
@@
 Thus, adapting the continuous learning techniques to handle such constraints is a critical challenge in the field of ML for high rate streaming of scientific data.
-The other challenge is to effectively navigate the memory hierarchy of heterogeneous computing systems, as it exposes enormous differences in throughput [12, 13].
+A challenge to scaling is to effectively navigate the memory hierarchy of heterogeneous computing systems, as it exposes enormous differences in throughput [12, 13].
 While internal throughput in GPUs can easily reach the TB/s range on a single node, breaking down the throughput of massively parallel filesystems to the single node throughput reveals maximum performance of several tens of MB/s only.
 Each node often houses more than one GPU, hence available data storage and bandwidth demand data reduction, often done by discarding highly valuable data in practice.
-We further note that the mismatch in throughput between memory hierarchies in modern HPC systems has been steadily growing, increasing the number of cases in which offline analysis is impossible and thus the urgency to find solutions for online analysis of data at high rates and volumes [12].
+We further note that mismatch in throughputs between memory hierarchies in modern HPC systems has been steadily growing, increasing the number of cases in which offline analysis is impossible and thus the urgency to find solutions for online analysis of data at high rates and volumes [12, 14].
 We argue that in case of such extreme rates online analysis of streaming data is the method of choice [15, 7], reducing the overall amount of data that needs to be stored while using computationally demanding algorithms to make optimum use of the data produced.
@@
 We furthermore emphasize that besides presenting a full workflow, the additional unique value of this work is to highlight the multi-parameter optimization problem of fitting the available compute and data transfer resources to achieve the best result in the shortest amount of time.
-We remark that our technique is especially interesting in the case of irreversible data loss [16], and thus extends well beyond the case of large-scale simulations as high rate data sources presented here.
+We remark that our technique is especially interesting in the case of irreversible data loss [17].
+It thus extends well beyond the extreme case of large-scale simulations presented here.
+Examples include other high data rate sources, e.
+g.
+plasma or particle experiments with high-resolution and high-repetition-rate detectors, which are more likely connected to smaller edge clusters.
 With a goal to automatically extract knowledge from large scale simulations, the paper makes the following contributions:
@@
 This orchestration, to the best of our knowledge, is the first of its kind.
-Scientifically: We uncover correlations between emitted radiation and particle dynamics within the simulation in an unsupervised manner to create an inverse map from the observed radiation to the particle dynamics and to identify the aspects of the particle dynamics that are relevant to the remotely observable radiation signatures of KHI.
+Scientifically: We uncover correlations between emitted radiation and particle dynamics within the simulation in an unsupervised manner to create an inverse map from the observed radiation to the particle dynamics, and to identify the aspects of the particle dynamics that are relevant to the remotely observable radiation signatures of KHI.
 Using lessons learned from this orchestration and challenges faced, we provide input on what needs to be improved as we attempt to understand complex scientific data generated by instruments.
@@
 The Kelvin-Helmholtz instability (KHI) is a well-known shear surface instability observed in plasmas [8].
-It is driven by a self-amplifying cycle of small density or velocity fluctuations that lead to a growing magnetic field at the shear surface, which further amplifies the initial fluctuations as depicted in Fig.
-1.
+It is driven by a self-amplifying cycle of small density or velocity fluctuations that lead to a growing magnetic field at the shear surface, which further amplifies the initial fluctuations (as depicted in Fig.
+1).
 This phenomenon occurs in astrophysical scenarios, including Saturn’s magnetopause [19], but also in fusion reactors and inertial confinement fusion (ICF) [20, 21].
-While the theoretical understanding and simulation capabilities of KHI [21, 22] are well established, direct measurement of the predicted density changes remains a real challenge.
+While the theoretical understanding and simulation capabilities of KHI [22, 23] are well established, a direct measurement of the predicted density changes remains a real challenge.
 Because of its well-understood dynamics, we chose KHI as a test case for our study to validate the results of our orchestrated workflow against existing knowledge.
@@
 However, this correlation process typically requires extensive post-processing and analysis, often spanning several years.
-Here, we simulate the KHI using the plasma simulation code PIConGPU [24, 6] which utilizes the alpaka library [25] for performance portability and has been at the forefront of developing central technologies for GPU-accelerated plasma simulations.
+Here, we simulate the KHI using the plasma simulation code PIConGPU [25, 6] which has been at the forefront of developing central technologies for GPU-accelerated plasma simulations, utilizing the alpaka library [26] for performance portability.
 It was the first particle-in-cell (PIC) [27] code to run performant simulations on a variety of architectures in the world over many years [7].
@@
 Fig.
-2 depicts the interrelations between observational data, such as radiation, and simulation state, such as particles, as both inputs and outputs to data-driven models.
+2 depicts the interrelations between observational data (such as radiation) and simulation state (such as particles) as both inputs and outputs to data-driven models.
 Our objective of approximating the inverse mapping from observed radiation to particle dynamics can abstractly be described by the information flow depicted in Fig.
@@
 2(b).
-To observe the extent to which this representation captures relevant features of the particle dynamics, it is worthwhile to train a surrogate model of the, computationally expensive, forward process of radiation emission (Fig.
+To observe the extent to which this representation captures relevant features of the particle dynamics, it is worthwhile to train a surrogate model of the computationally expensive, forward process of radiation emission (Fig.
 2(c)).
@@
 It was shown in 2013 [6] that the computation of high quality, high resolution observational data such as radiation spectra can only be performed in-situ, as storing the ground truth data, the phase space trajectory data of each particle in the simulation at each time step, is impossible due to the aforementioned throughput hierarchy.
-As an example, scaling to a moderate 25 % of the Frontier system, one would be faced with 1 PB of data for every time step.
-With total time steps per simulation of the order of a few thousand time steps and time step durations of the order of 0.1 to 1 s, we frequently observe data rates of 1 to 10 PB/s for particle trajectory data and theoretically require total volumes on disk of the order of 10 EB.
-Depleting storage in as little as .
-We thus end up with the conundrum to employ data-driven models for solving a large-scale ill-posed inverse problem.
+As an example, scaling to even a moderate 25 % of the Frontier system, one would be faced with 1 PB of data for every time step.
+With total time steps per simulation of the order of a few thousand time steps and time step durations of the order of 0.1 to , we frequently observe data rates of 1 to 10 PB/s for particle trajectory data and theoretically require total volumes on disk of the order of 10 EB.
+This means the storage could get depleted for a simulation that is as short as in duration.
+We thus end up with the conundrum to employ data-driven models for solving a large-scale ill-posed inverse problem without sufficient memory.
 III-A A complex in-memory workflow to reconstruct local phase space dynamics from radiation
-A particular challenge for the reconstruction of local phase space dynamics is that their radiation signatures are highly convoluted, composite quantities that combine multiple facets of the ground truth which to restore.
-This observation guides our approach at learning from a live PIC simulation, as the posed problem is now diagnosed as ill-posed.
+A particular challenge for the reconstruction of local phase space dynamics is that their radiation signatures are highly convoluted, composite quantities that combine multiple facets of the ground truth to be restored.
+This observation guides our approach at learning from a live PIC simulation, as the posed problem is now diagnosed as ill-posed [28].
 In detail, the observed radiation spectrum is both spectrally and angularly resolved and the observed intensity at each wavelength and direction varies between time steps according to the changes of the phase space dynamics.
 The observed intensity in each direction and at each wavelength is a weighted sum of all contributions from all particles.
-Hence here are the steps to be taken:
-Simulate a sufficiently large domain at high spatial and temporal resolution to capture both the local plasma dynamics as well as the collective contributions of different plasma regions to the total radiation spectrum.
+Hence, here are the steps to be taken:
+Simulate a sufficiently large domain at a high spatial and temporal resolution to capture both the local plasma dynamics, as well as the collective contributions of different plasma regions to the total radiation spectrum.
 At each PIConGPU time step:
-Calculate the radiation spectrum directly in-situ as the sum of the contributions from all radiating particles in the simulation at each time step, with high angular and spectral resolution and bandwidth.
-Collect the particle phase space data, positions and momenta, and the complex amplitude and phase for each wavelength and direction of the radiation.
-Prepare the collected data for an ML model by finding a suitable encodings for spectral and phase space data.
-Stream that data to an ML model and asynchronously train the model.
-Stop the simulation after observing the complete physical process under investigation.
+Calculate the radiation spectrum in-situ as the sum of the contributions from all radiating particles in the simulation at each time step, with high angular and spectral resolution and high bandwidth.
+Collect the particle phase space data (namely the positions and momenta) and the complex amplitude and phase for each wavelength and direction of the radiation.
+Prepare the collected data for an ML model by finding suitable encodings for spectral and phase space data.
+Stream the encoded data to an ML model and asynchronously train the model.
+Stop the simulation after observing and covering a sufficient amount of relevant stages of plasma instability simulation.
 III-B The technical challenge: Exascale data streaming and continual learning
 An important design goal for the implementation of the above data processing scheme is to write no simulation data to the filesystem during the entire workflow.
-File I/O can certainly be initiated when desired, but is independent of and not necessary for the workflow.
-While in-situ approaches have been previously used to compute scientifically relevant analyses through plugins within the same application context [6], this is no longer possible for the envisioned ML analysis, as the vastly incompatible software stacks for the C++ simulation code PIConGPU and the PyTorch-based MLapp as well as their different scaling properties inhibit such an integration into a strongly-coupled monolithic application framework.
-Hence, simulation and the MLapp are modularized into two separate applications that communicate through a standardized data interface, a pattern that we refer to as loose coupling, since the standardized data interface allows to plug together a data processing pipeline more flexibly, i.e.
+File I/O can certainly be initiated when desired, but is not necessary for the workflow.
+In-situ approaches have been previously used to compute scientifically relevant analyses through plugins within the same application context [6].
+This is no longer possible for the envisioned ML analysis, as the vastly incompatible software stacks for the C++ simulation code PIConGPU and the PyTorch-based MLapp, as well as their different scaling properties inhibit such an integration into a strongly-coupled monolithic application framework.
+Hence, the simulation and the MLapp are modularized into two separate applications that communicate through a standardized data interface – a pattern that we refer to as loose coupling, since the standardized data interface allows to plug together a data processing pipeline more flexibly, i.e.
 loosely than before.
@@
 The latter becomes necessary when transitioning from file-based workflows to in-memory workflows without sacrificing flexibility in the data exchange pattern, as well as in response to codes that scale differently.
-The whole simulation, all data produced by the simulation and ingested by the model, and the whole model all reside in-memory and data is distributed and optionally reduced in-transit (Fig. 3(b)).
+The whole simulation, all data produced by the simulation and ingested by the model, and the whole model, all reside in-memory and data is distributed and optionally reduced in-transit (Fig. 3(b)).
 The physical workflow outlined in III-A can be directly translated into a technical setup that is the main subject of this work:
-Start a scalable, GPU-accelerated PIConGPU simulation of sufficient size and resolution, utilizing a large subset of nodes on Frontier.
-Schedule PyTorch and N/RCCL alongside the PIConGPU simulation. (N/RCCL - NVIDIA and ROCm collective communication library)
+Start a scalable, GPU-accelerated PIConGPU simulation of sufficient size and resolution, utilizing a large subset of nodes on Frontier [29].
+Schedule PyTorch and N/RCCL (i. e. NVIDIA and ROCm collective communication library) alongside the PIConGPU simulation.
 Optimize scheduling for data transport and locality, physical system size and resolution, simulation and training time, total time of solution and use of resources (options in Fig. 3(c)).
-At each simulation time step
-stream particle momenta and positions and spectral data to a large-scale ML model,
+At each simulation time step:
+Stream particle momenta, particle positions and the spectral data to a large-scale ML model,
 Transform the phase space and spectral data from its simulation format to an optimum format for the ML mode and
@@
 Repeat the workflow for a sufficient number of time steps to cover all relevant stages of the plasma instability simulation with respect to the dynamical evolution and amount of training data needed.
-For supporting an online workflow which stores no intermediate data such as phase space trajectory and radiation data to disk, the model must be continuously trained online from the series of time step data.
+For supporting an online workflow which stores no intermediate data such as phase space trajectory and radiation data to disk, the model must be continuously trained online from the series of time step.
 This requires the model to be trained by subsequent snapshots of the evolution of plasma and radiation data.
 For this, continual learning is the methodology of choice as outlined in Section IV-C.
-We strongly believe that a large class of inversion problems for complex, large-scale systems that produce data streams face similar challenges and that our solution presented here for the first time introduces all relevant building blocks and technologies to address these challenges.
+We strongly believe that a large class of inversion problems for complex, large-scale systems that produce data streams face similar challenges, and that our solution presented here for the first time introduces all the relevant building blocks and technologies to address these challenges.
 IV Methodology
@@
 the weighted sum of the total number of particle updates per second (90 %) and the number of cell updates per second (10 %), is demonstrated as shown in the weak scaling results.
-In this particular configuration, the simulation included macroparticles in grid cells, compared to our base FOM run on Summit in 2019, where only macroparticles in cells could be used.
+For the results presented in this paper, the simulation involved macroparticles in grid cells, compared to our base FOM run on Summit in 2019, where only macroparticles in cells could be used [31].
 On Frontier, one thousand time-steps completed in a mere 6.5 minutes.
@@
 IV-B Data streaming via openPMD
-Considering that PIConGPU as well as the MLapp both rely significantly on very different software stacks, launching both parts of the setup in the same application context is not feasible.
+Considering that both PIConGPU as well as the MLapp rely significantly on very different software stacks, launching both parts of the setup in the same application context is not feasible.
 As detailed previously in III-B, we follow a loose-coupling rather than a tight-coupling approach [15], launching both applications separately.
@@
 the provisioned Orion filesystem, cannot be expected to scale up to full system size as a base for a loosely coupled data analysis pipeline.
-The in-transit data strategy is built upon the Sustainable Staging Transport (SST) [39, 40] data engine of ADIOS2 that connects one parallel data producer code to an arbitrary number of parallel data consumer codes, hereby opening connections only between processes that actually share data.
+The in-transit data strategy is built upon the Sustainable Staging Transport (SST) [42, 43] data engine of ADIOS2 that connects one parallel data producer code to an arbitrary number of parallel data consumer codes, thereby opening connections only between processes that actually share data.
 For adapting to different system architectures, the SST engine implements different network transport technologies (data planes), including TCP (non-scalable fallback), libfabric, ucx and the MPI_Open_port() API of MPI.
@@
 At each scaling run (weak scaling), five time steps are sent from PIConGPU to the no-op consumer, which then measures the time needed for loading the data.
-The perceived parallel throughput is computed based on this measured time and the global data size, i.
+The parallel throughput is calculated based on this measured time and the global data size, i.
 e.
 the previously mentioned per node multiplied with the number of nodes.
-While including communication overhead, this perceived parallel throughput has been shown in [40] to be a close approximation for the real throughput in this kind of setup.
+Although including communication overhead, this throughput value has been shown in [43] to be a close approximation for the real throughput in this kind of setup.
 The benchmarks in Fig.
@@
 Depicted are boxplots of all single measurements.
-Most notably, we observe a maximum perceived parallel throughput of 20 - 30 TB/s which compares outstandingly against the 10 TB/s bandwidth of the parallel Orion filesystem, a scaling limit which we circumvent and exceed by not using the filesystem for intermediate data.
+Most notably, we observe a maximum parallel throughput of 20 - 30 TB/s which compares outstandingly against the 10 TB/s bandwidth of the parallel Orion filesystem, a scaling limit which we circumvent and exceed by not using the filesystem for intermediate data.
 Furthermore, these results can also compare well against the 35 TB/s aggregate write bandwidth of the SSDs installed locally in the compute nodes [45].
@@
 6).
-While these results signify the best achieved per-node throughput of 3.5 - 4.7 GB/s at 4096 compute nodes (a total of 14.1 - 18.8 TB/s), it turned out that this strategy did not scale to the full system, so we employed an alternative that enqueued operations in batches of 10 operations.
+While these results signify the best achieved per-node throughput of 3.5 - 4.7 GB/s at 4096 compute nodes (a total of 14.1 - 18.8 TB/s), it turned out that this strategy did not scale to the full system.
+Thus, we employed an alternative that enqueued operations in batches of 10 operations.
 While this turned out to scale to the full system, it came at a notable cost in performance, visible in a per-node throughput of 1.9 - 2.6 GB/s at 9126 compute nodes (a total of 16.5 - 23.0 TB/s).
@@
 The task of in-transit ML is to extract patterns from the live observations and thereby learn correlations between variables and reduced representations.
-Here, we focus on predicting particle dynamics from radiation spectra, which constitutes an ill-posed problem in that any given radiation spectrum can potentially be produced by various types of dynamics.
+Here, we focus on predicting particle dynamics from radiation spectra, which constitutes an ill-posed problem – since any given radiation spectrum can potentially be produced by various types of dynamics [28].
 When learning in-transit from a simulation, this problem is reduced from an intractable general case to the question of: which type of local particle dynamics under the given physical setup, i.
 e.
-the relativistic KHI, gave rise to an observed radiation spectrum observed in one direction, which, however, remains an ill-posed problem.
+the relativistic KHI, gave rise to an observed radiation spectrum observed in one direction, which, however, remains an ill-posed problem [28].
 Many ML techniques, especially feed-forward networks, tend to predict the mean of possible inversions, even if this itself is not a valid solution [46].
 One class of techniques specifically designed to learn by sampling from conditional multi-modal distributions, such as those that occur as solutions to our ill-posed problem, are flow-based models [47, 48] and invertible neural networks (INNs) [49].
-These learn an invertible transformation, or change-of-variables, between a chosen posterior distribution, often a normal distribution , and a prior described by the training data, in our case the distribution of observed particle dynamics .
-One specific difference between them is, that INNs also learn the deterministic forward function defining the inverse problem, in our case the mapping between samples of particle dynamics and radiation spectrum .
-Put differently, they also learn to predict the condition under which a given sample of the prior may occur.
+These learn an invertible transformation from a chosen posterior distribution (often a normal distribution ) and a prior described by the training data (in our case the distribution of observed particle dynamics ) and vice versa.
+One specific difference between them is that INNs also learn the deterministic forward function defining the inverse problem – in our case, the mapping between samples of particle dynamics and radiation spectrum .
+Put differently, INNs also learn to predict the condition for a given prior .
 Particle configurations in phase space of the plasma dynamics simulated here can be very complex, such that a small sample of particles cannot adequately represent the dynamics.
-For INNs and flow-based models this poses a challenge, because the change-of-variables approach [47, 48] dictates, that the information volume, i.
+For INNs and flow-based models this poses a challenge, because the change-of-variables approach [50, 51] dictates that the information volume, i.
 e.
 the number of activations at the input and output of each coupling block, must be constant throughout the network, which results in a very large network when working with a prior distribution of dimensionality on the order of .
-In this setting, the number of network parameters can be kept in check by using affine flows [49, 44] with designed convolutional sub-networks which should also be invariant to transposition of particles in the input vector to be efficiently trainable.
+In this setting, the number of network parameters can be kept in check by using affine flows [52, 47] with convolutional sub-networks designed to be invariant to transposition of particles in the input vector to be efficiently trainable.
 When generating point cloud (PC), flow-based models are typically tasked with sampling them point-by-point, i.e.
-the model learns a probability density that describes the particle density in phase space.
-The main drawback here is, that the model has to learn the many-modal distributions which can occur in the KHI phase space on a single-particle basis and can not longer generate multiple particle distribution for a single condition, i.e.
-not distinctly produce all PCs solving an ill-posed problem.
-Another issue is, that, unless we use an encoder (conditioning network) producing very small latent vectors of around six dimensions for the radiation, the dimension of the condition is much larger than the dimension of the prior while the network has to be sampled times to retrieve a useful particle distribution.
-To overcome the above inefficiencies, we opted to view the particle configuration as a point cloud (PC) and reduce the dimensionality of the prior by tasking an outer network to learn latent representations of the phase space, which are also invariant with respect to transposition of particles.
-This outer network can in principle be an autoencoder (AE) whose latent representation is passed to the INN as the prior .
+the model learns a probability density describing the particle density in phase space.
+The main drawback of this method is that the model has to learn the many-modal distributions which can occur in the KHI phase space on a single-particle basis and can no longer generate multiple particle distribution for a single condition, i.
+e.
+we cannot expect it to distinctly produce all PCs solving an ill-posed problem [28].
+Another issue is that, unless we use an encoder (conditioning network) producing very small latent vectors of around six dimensions for the radiation, the dimension of the condition is much larger than the dimension of the prior while the network has to be sampled times to retrieve a useful particle distribution.
+To overcome the above inefficiencies, we opted to 1.
+view the particle configuration as a point cloud (PC), and 2.
+reduce the dimensionality of the prior by tasking an outer network to learn latent representations of the phase space that are also invariant with respect to transposition of particles.
+This outer network can, in principle, be an autoencoder (AE) whose latent representation is passed to the INN as samples of the prior distribution .
 In practice, the INN will not generate exact latent vectors on its backward pass, but variations, which will cause the AE’s decoder to fail at reconstruction.
@@
 As depicted in Fig. 7, there are three main blocks of the network architecture:
-The encoder (light green) follows an architecture proposed in [50], which by itself is a simplified version of PointNet [51], using convolutions acting on the coordinates of particles separately to gradually (channels: ) extract 608 features per particle, followed by a max-pooling to obtain a transposition-invariant feature set, which is then processed by two multi-layer perceptrons (MLP), each with hidden features, to obtain predictions for the mean and standard deviation to sample the 544-dimensional latent vector .
+The encoder (light green) follows an architecture proposed in [53], which is a simplified version of PointNet [54].
+6-dimensional vectors corresponding to the positions and momenta of the particles are fed to convolutions separately to extract 608 features per particle (channels: ), followed by a max-pooling to obtain a transposition-invariant feature set.
+These 608 features are then processed by two multi-layer perceptrons (MLP), each with hidden features, to obtain predictions for the mean and standard deviation to sample the 544-dimensional latent vector .
 The decoder (cyan) uses a single fully-connected layer (, reshaped to to transform the latent before using 3D deconvolutions (channels: , kernelsize: , stride ) to upsample to particles.
@@
 The loss function to be minimized during training of this architecture consists of five terms, two connected to the VAE and three for the INN, detailed below:
-As the reconstruction loss of the VAE, we use Chamfer’s distance (CD) [53], which is the most commonly used differentiable function to compare PCs in literature, because it is efficient to compute.
-Regularization of the VAE’s latent space is controlled by Kullback-Leibler divergence (KL) [54].
-The resulting spectrum of the INNs forward function is compared to the ground truth using mean-squared error
-The posterior normal distribution, produced by the INN’s forward pass, and the prior distribution, produced by the INN’s backward pass, are both evaluated using maximum mean discrepancy (MMD) using an inverse multi-quadratic kernel, as proposed in [46].
-The total loss corresponds to the weighted sum
+: As the reconstruction loss of the VAE, we use the Chamfer’s distance (CD) [56], it being the most popular differentiable function for comparing PCs the in literature, owing to it being cheap to compute.
+: Regularization of the VAE’s latent space is controlled by Kullback-Leibler divergence (KL) [57].
+: The resulting spectrum of the INNs forward function is compared to the ground truth using mean-squared error
+, : Both the posterior normal distribution produced by the INN’s forward pass and the prior distribution produced by the INN’s backward pass are evaluated using maximum mean discrepancy (MMD) using an inverse multi-quadratic kernel [49].
+The total loss corresponds to the weighted sum , where,
 The numerical values for the coefficients were empirically tuned to optimize performance of the AE reconstruction and the INN forward and backward predictions.
@@
 To integrate with the streaming setup, we implemented this approach to continual learning as a separate component, called training buffer, between the streaming receiver and the training loop.
-This component holds two buffers: a now-buffer, which holds the latest samples received.
+This component holds two buffers: a now-buffer, and an EP-buffer.
+The now-buffer holds the latest samples received.
 Each time new samples are received, they are prepended to the now-buffer and any element which would then exceed the buffer size is removed and added to the EP buffer.
@@
 This is crucial to allow the optimizer some amount of exploration, which can only happen sequentially and hence a smaller number of training iterations cannot be compensated by the large batch sizes of data-parallel training.
-For all training runs, we use the Adam optimizer [56] with , , and an weight decay strength of .
+For all training runs, we use the Adam optimizer [59] with , , and weight decay .
 For data parallel training we employ PyTorch’s Distributed Data Parallel (DDP) module with its N/RCCL backend.
@@
 It also fails to distinguish some point configurations [61].
-Both of these problems are overcome by the earth mover’s distance (EMD) [59] loss, but at much higher computational cost111We observed about a increase in batch times when using EMD [60] vs.
-a simple implementation of CD..
+Both of these problems are overcome by the earth mover’s distance (EMD) [62] loss, but at much higher computational cost111We observed about a increase in batch run times when using EMD [63] compared to a simple implementation of CD..
 Still, a rather compute and memory efficient differentiable implementation of EMD is available in the geomloss python package [63], which, instead of relying entirely on PyTorch, internally uses the KeOps library [64, 65] for performance.
@@
 IV-D Setup on ORNL Frontier system
-While distributing nodes exclusively to either PIConGPU or to the MLapp, inter-node setup in Fig.
-3(c), is easier to achieve in Slurm (and this strategy would be supported by our flexible I/O solution), we decide to use the intra-node setup in Fig.
-3(c) and share each compute node between PIConGPU and the MLapp.
-This has the advantage that data exchange mostly needs not leave the node, but requires a more specific resource allocation: Each single compute node no longer runs homogeneous parallel sub-jobs, but has to locally assign resources to heterogeneous applications.
-For our setup, this local assignment gives 4 GCDs to PIConGPU and 4 GCDs to MLapp.
+While distributing nodes exclusively to either PIConGPU or to the MLapp is easier to achieve in Slurm (i.
+e.
+inter-node setup in Fig.
+3(c) that is supported by our flexible I/O solution), we decide to use the intra-node setup from Fig.
+3(c) where each compute node is shared between PIConGPU and the MLapp.
+This strategy has the advantage that data exchange mostly does not need to leave the node, but requires a more specific resource allocation.
+More specifically, each single compute node no longer runs homogeneous parallel sub-jobs, but has to locally assign resources to heterogeneous applications.
+For our setup, this local assignment gives 4 Graphics Compute Dies (GCDs) to PIConGPU and 4 GCDs to MLapp.
 In openPMD and ADIOS2, each reader application decides on its own which remote datasets to load.
@@
 Due to the plugin-based structure of PIConGPU, both the particle and radiation input required by MLapp are provided by distinct output plugins of PIConGPU, meaning that two parallel data streams are opened between PIConGPU and the MLapp.
-Each plugin prepares the current time step’s data and, upon having all data ready, ADIOS2 will gather the metadata, including remote read addresses, to parallel rank 0 and provide the time step to the reader.
+Each plugin prepares the current time step’s data and, upon having all data ready, ADIOS2 will gather the metadata, including remote read addresses to parallel rank 0 and provide the time step to the reader.
 The reader performs remote reads to its liking and closes the time step, indicating to the writer that the data can now be dropped.
 A unique and important requirement of the presented workflow is the ability to not only run but also scale on a large number of Frontier’s AMD MI250X GPUs.
-Our machine learning model is small enough to fit on a single GCD; therefore, parallel training of this model is done using data parallelism where copies of the model are distributed across GCDs with each copy of the model receiving different chunks of data to train on.
+As our machine learning model is small enough to fit on a single GCD, parallel training of this model is done using data parallelism, where copies of the model are distributed across GCDs with each copy of the model receiving different chunks of data to train on.
 Once each model computes its gradients, all the instances of the model must do a collective all-reduce communication to average the gradients.
@@
 We define efficiency as the ratio of runtime at a specific size over runtime of the smallest size.
-Before averaging the times per training batch, we removed outliers more than four standard-deviations from the mean, because on Frontier we observed single batches taking more than the mean time.
+Before averaging the times per training batch, we removed outliers more than four standard-deviations from the mean, as we observed single batches taking more than the mean time on Frontier.
 While ideal scaling should correspond to 100 % efficiency across all sizes, communication overheads reduce efficiency in practice.
 Here, the efficiency of in-transit training reaches around 35 % at 96 nodes.
-We attribute the low parallel efficiency to two main reasons: One of them is the inevitable all-to-all communication between PyTorch ranks that takes place to average the gradients during each backward pass and obtain a synchronized result on all ranks, which accounts for a deficit of , which is still a respectable result achieved by overlapping computations and communication in PyTorch DDP.
-The other is the computation of the two MMD loss terms, which amount to matrix dot products with data distributed across all ranks.
-Due to torch not offering a distributed primitive and the matrices involved being relatively small, even at the largest achieved scale, we opted for a naive implementation which replicates work between nodes.
+We attribute the low parallel efficiency to two main reasons.
+The first reason is the inevitable all-to-all communication between PyTorch ranks that takes place to average the gradients during each backward pass and obtain a synchronized result on all ranks, which accounts for a deficit of , which is still a respectable result achieved by overlapping computations and communication in PyTorch DDP.
+The second reason is the computation of the two MMD loss terms, which amount to matrix dot products with data distributed across all ranks.
+As torch (less than version 2.2) did not offer a distributed primitive and the matrices involved are relatively small, even at the largest achieved scale, we opted for a naive implementation which replicates work between nodes.
 The global communication required here222torch.distributed.all_gather_into_tensor breaks the torch computational graph, i.e.
@@
 That is, the increased learning rate overcompensates the training slow down due to larger batch sizes.
-Furthermore, it became evident that the three losses connected to the INN converged best at lower learning rates, while the VAE only found better minima at the highest learning rate, showing that separate learning rates and of the encoder/decoder block and the inversion block, respectively, need to be applied at large scales.
-Within the scope of this study, we have to conclude, that for in-transit training at very large batch sizes, hyper parameter studies do not transfer from small-scale experiments and thus have to be performed at scale.
+Furthermore, it became evident that the three losses connected to the INN converged best at lower learning rates, while the VAE only found better minima at the highest learning rate, showing that separate learning rates and for the encoder/decoder block and the inversion block respectively need to be applied at large scales.
+Within the scope of this study, we have to conclude that for in-transit training at very large batch sizes, hyper parameter studies do not transfer from small-scale experiments and thus have to be performed at scale.
 To enable in-transit learning at the scales we tested or beyond, comprehensive studies of the relations between the block learning rates and , batch sizes, and maybe even loss weights in Eq.
@@
 Exploring this parameter space for the large scale training requires the full study to be systematically performed at these scales (up to 384 GCDs), either by employing a streamed simulation or the same training data stored on disk.
-We argue, that the latter is challenging at scale due to the lack of disk capacity and bandwidth.
+We argue that the latter is challenging at scale due to the lack of disk capacity and bandwidth.
+Efficient approaches to autotuning and optimization will be indispensable for this next step.
 V-B Quantifying the predictive capabilities of the machine learning method
@@
 The ML reconstruction consistently predicts these two distinct populations.
-This is consistent with the interpretation, that the INN has learned unsupervised that the vortex region undergoes more acceleration and thus emits more intense radiation than the bulk of the plasma.
+This is consistent with the interpretation that the INN has learned unsupervised that the vortex region undergoes more acceleration and thus emits more intense radiation than the bulk of the plasma.
 However, the momentum prediction is less accurate: the mean of (one of) these populations often does not match the PIConGPU data (Fig.
@@
 We suspect that this is partly due to our choice of region boundaries on the shear surface, which results in a much lower sampling rate of one flow direction compared to the other, leading to a lower accuracy in predicting the mean of the lower peak of the momentum distribution.
-It is also possible, that this problem arises from the VAE’s decoder having difficulties reconstructing discontinuous distributions.
+It is also possible that this problem arises from the VAE’s decoder having difficulties reconstructing discontinuous distributions.
 Nevertheless, the successful prediction of shear regions indicates that the encoder extracts crucial features from the input data to recognize whether a region contains a vortex and encodes them in latent space.
-Despite suboptimal reconstruction of vortex momenta, the ML prediction still allows to clearly identify KHI regions.
+The model clearly learned to partition the latent space into regions for different flow directions and vortex regions, which both the encoder and the inversion network learned to map to.
+These features are extracted using loss functions that do not impose physical rules and still encode physical properties of the system.
+Hence, in principle, they allow a simple, almost linear classifier to predict physical regimes.
+Evaluating such a classifier quantifies successful extraction of underlying physical rules.
+We observed that main issue of our model is suboptimal reconstructions of vortex momenta, which we attribute to shortcomings of our decoder architecture, despite which the ML prediction still allows to visually identify KHI regions.
 Previous work on PC generation [67] observed that the latent space of an AE may contain more information than the decoder can reconstruct with satisfactory fidelity and that more sophisticated generative models, such as GANs or flows, can learn to produce more detailed reconstructions when conditioned on the same latent space.
@@
 This allows us to circumvent the lack of adequate disk capacity and bandwidth, often a bottleneck in large-scale systems, by enabling data to reside in-memory and be distributed via network between the nodes.
-Scientifically, this work also demonstrates as a proof of principle, that it is possible to learn correlations from a physics simulation on-the-fly.
+Scientifically, as a proof of principle, this work also demonstrates that it is possible to learn correlations from a physics simulation on-the-fly.
+The presented workflow is scalable from local clusters to Top-1 supercomputers and is tested across GPU vendors.
+This is possible due to PIConGPU’s foundation on the abstraction layer alpaka and due to the availability of various ports of PyTorch.
 On full Frontier, we demonstrate streaming at scale using openPMD as well as performance studies at scale for general plasma physics.
 Acknowledgments
+The authors thank Patrick Stiller and Nico Hoffmann for fruitful discussions.
 This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, supported by the Office of Science of the U.S.
@@
 DE-AC05-00OR22725.
-The research was supported by Laserlab-Europe EU-H2020 GA no. 871124.
+This research was supported by the Helmholtz Association Initiative and Networking Fund in the frame of Helmholtz AI, by Laserlab-Europe EU-H2020 GA no.
+871124 and the European Union through Grant Agreement No.
+101093261 (Plasma-PEPSC).
 The support work from the ADIOS team was supported by the U.S.